Brown sold his car for 3/5 of the cost price and finds that in doing so he has list £750 .How much did he sell the car?
Answer by Theo(13342)


if he sells the car for 750 and 750 was 3/5 of the cost, then the cost of the car would be 5/3 * 750 = 1250.

you would solve that as follows:

let x = the cost of the car.

3/5 * x = 750.

multiplhy both sides of this equation by 5/3 to get x = 5/3 * 750.

this makes x = 1250, which is the cost of the car.

if the cost of the car is 750 and he sells it for 3/5 of the cost, then he sells is for 3/5 * 750 = 450.

in that case, you would solve it as follows.

let x = the price he sold the car for.

than x = 3/5 * 750.

solve for x to get x = 450.